OpenStudy (anonymous):

Find the domain of each function a. f(x)=x/x^2-9 b. g(x)=3/sqrt of x+7 c. h(x)=sqrt of 5-x

OpenStudy (anonymous):

is the first one \(f(x)=\frac{x}{x^2-9}\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so, domain is all real numbers except where the denominator is zero set \(x^2-9=0\) and solve for \(x\)

OpenStudy (anonymous):

why

OpenStudy (anonymous):

for the next two, make sure that a) \(x+7>0\) and b) \(5-x\geq 0\)

OpenStudy (anonymous):

oh, to answer your question, because you cannot divide by \(0\)

OpenStudy (anonymous):

so if you write \(\frac{-3}{0}\) it is not a number

OpenStudy (firejay5):

@satellite73 Fan me please

OpenStudy (anonymous):

Ok, thnx for the useful info

OpenStudy (anonymous):

yw